square root of (3x+3)+ square root of (x+2)= 5, solve for x. The answer is 2, need the steps for the answer.
Answer by edjones(8007)

sqrt(3x+3)+sqrt(x+2)=5
sqrt(3x+3)=5-sqrt(x+2)
3x+3=x+27-10sqrt(x+2) Square each side.
2x-24=-10sqrt(x+2)
4x^2-96x+576=100x+200 square each side.
4x^2-196x+376=0
4(x^2-49x+94)=0
(x-47)(x-2)=0
x=2 is the only answer. x=47 is extraneous.
.
Ed